Coffee Bean Inc. (CBI) processes and distributes a variety of coffee. CBI buys coffee beans from around the world and roasts, blends, and packages them for resale. Currently, the firm offers 15 coffees to gourmet shops in 1-pound bags. The major cost is direct materials; however, a substantial amount of factory overhead is incurred in the predominantly automated roasting and packing process. The company uses relatively little direct labor.

Some of the coffees are very popular and sell in large volumes; a few of the newer brands have very low volumes. CBI prices its coffee at full product cost, including allocated overhead, plus a markup of 30%. If its prices for certain coffees are significantly higher than the market, CBI lowers its prices. The company competes primarily on the quality of its products, but customers are price conscious as well.

Data for the current budget include factory overhead of $3,100,000, which has been allocated on the basis of each product’s direct labor cost. The budgeted direct labor cost for the current year totals $601,000. The firm budgeted $6,100,000 for purchase and use of direct materials (mostly coffee beans).

The budgeted direct costs for 1-pound bags of two of the company’s many products are as follows:


Mona LoaMalaysian
Direct materials$4.20
$3.20
Direct labor
0.30

0.30

CBI’s controller, Mona Clin, believes that its current product costing system could be providing misleading cost information. She has developed this analysis of the current year’s budgeted factory overhead costs:

ActivityCost DriverBudgeted ActivityBudgeted Cost
PurchasingPurchase orders
1,168
$580,000
Materials handlingSetups
1,810

721,000
Quality controlBatches
730

145,000
RoastingRoasting hours
96,200

962,000
BlendingBlending hours
33,700

337,000
PackagingPackaging hours
26,100

261,000
Total factory overhead cost



$3,006,000

Data regarding the current year’s production of just two of its lines, Mona Loa and Malaysian, follow. There is no beginning or ending direct materials inventory for either of these coffees.


Mona LoaMalaysian
Budgeted sales101,000 pounds2,010 pounds
Batch size10,100 pounds510 pounds
Setups3 per batch3 per batch
Purchase order size25,100 pounds510 pounds
Roasting time1 hour per 100 pounds1 hour per 100 pounds
Blending time0.5 hour per 100 pounds0.5 hour per 100 pounds
Packaging time0.1 hour per 100 pounds0.1 hour per 100 pounds

Required:

1. Using Coffee Bean Inc.’s current product costing system,

a. Determine the company’s predetermined overhead rate using direct labor cost as the single cost driver.

b. Determine the full product costs and selling prices of one pound of Mona Loa coffee and one pound of Malaysian coffee.

2. Using an activity-based costing approach, develop a new product cost for 1 pound of Mona Loa coffee and 1 pound of Malaysian coffee. Allocate all overhead costs to the 101,000 pounds of Mona Loa and the 2,010 pounds of Malaysian.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Coffee Bean Inc.
Ans 1 a
Calculation of predetermined overhead rate Amount $ Note
Budgeted factory overhead            3,006,000.00 A
Budgeted direct labor cost               601,000.00 B
Predetermined overhead rate                           5.00 C=A/B
Ans 1 b- Workings Mona Loa Malaysian Note
Direct labor cost                           0.30                     0.30 D
Predetermined overhead rate                           5.00                     5.00 See C
Overhead allocated                           1.50                     1.50 E=C*D
Ans 1 b
Unit Product cost Mona Loa Malaysian Note
Direct Materials                           4.20                     3.20
Direct Labor                           0.30                     0.30 See D
Overhead allocated                           1.50                     1.50 See E
Cost per 1 pound bag                           6.00                     5.00 F
Markup at 30%                           1.80                     1.50 G=F*30%
Sell price of 1 pound bag                           7.80                     6.50 H=F+G
Ans 2- Workings Mona Loa Malaysian Note
Expected Sales               101,000.00             2,010.00 I
Batch Size                 10,100.00                 510.00 J
Number of Batches                         10.00                     3.94 K=I/J
Setups per Batch                           3.00                     3.00 L
Number of Setups                         30.00                   11.82 M=K*L
Purchase Orders Size                 25,100.00                 510.00 N
Purchase Orders                           4.02                     3.94 O=I/N
Roasting time per 100 pound                           1.00                     1.00 P
Roasting Hours                   1,010.00                   20.10 Q=I/100*P
Blending time per 100 pound                           0.50                     0.50 R
Blending Hours                       505.00                   10.05 S=I/100*R
Packaging time per 100 pound                           0.10                     0.10 T
Packaging Hours                       101.00                     2.01 U=I/100*T
See workings above.
Overhead allocation table V W X=V/W Y Z=X*Y AA=Z/107000 AB AC=X*AB AD=AC/2070
Mona Loa Malaysian
Activity Cost Driver Cost Pool Cost Driver Volume Cost per Cost Driver Cost Driver used Cost allocated Cost Per unit Cost Driver used Cost allocated Cost Per unit
Purchasing Purchase Orders         580,000.00                        1,168.00                            496.58                          4.02              1,998.17                      0.02                          3.94              1,957.09                      0.95
Material Handling Number of setups         721,000.00                        1,810.00                            398.34                        30.00            11,950.28                      0.11                        11.82              4,709.81                      2.28
Quality Control Number of batches         145,000.00                           730.00                            198.63                        10.00              1,986.30                      0.02                          3.94                 782.84                      0.38
Roasting* Roasting Hours         962,000.00                      96,200.00                              10.00                  1,010.00            10,100.00                      0.09                        20.10                 201.00                      0.10
Blending Blending Hours         337,000.00                      33,700.00                              10.00                     505.00              5,050.00                      0.05                        10.05                 100.50                      0.05
Packaging Packaging Hours         261,000.00                      26,100.00                              10.00                     101.00              1,010.00                      0.01                          2.01                    20.10                      0.01
Total Activity Cost     3,006,000.00            31,084.75                      0.29              7,751.24                      3.74
Expected Sales          101,000.00              2,010.00
Ans 2
Unit Product cost Mona Loa Malaysian
Direct unit costs Amount $ Amount $ Amount $ Amount $
Direct Materials                           4.20                                3.20
Direct Labor                           0.30                     4.50                                0.30                                 3.50
Indirect unit costs
Purchasing                     0.02                                 0.95
Material Handling                     0.11                                 2.28
Quality Control                     0.02                                 0.38
Roasting                     0.09                                 0.10
Blending                     0.05                                 0.05
Packaging                     0.01                                 0.01
Total Unit cost                     4.80                                 7.25
Answer #2

Screen Shot 2021-09-27 at 10.58.48 AM.png

answered by: Maggie Wright
Answer #3

Screen Shot 2021-09-27 at 10.58.38 AM.pngScreen Shot 2021-09-27 at 10.58.30 AM.png

answered by: Maggie Wright